# Dissecting Query-Key Interaction in Vision Transformers

### v2

@@ Abstract
Self-attention in vision transformers is often thought to perform perceptual grouping where tokens attend to other tokens with similar embeddings, which could correspond to semantically similar features of an object. However, attending to dissimilar tokens can be beneficial by providing contextual information. We propose to use the Singular Value Decomposition to dissect the query-key interaction (i.e. ${\textbf{W}_{q}}^{\top}\textbf{W}_{k}$). We find that early layers attend more to similar tokens, while late layers show increased attention to dissimilar tokens, providing evidence corresponding to perceptual grouping and contextualization, respectively. Many of these interactions between features represented by singular vectors are interpretable and semantic, such as attention between relevant objects, between parts of an object, or between the foreground and background. This offers a novel perspective on interpreting the attention mechanism, which contributes to understanding how transformer models utilize context and salient features when processing images.
Figure 1: We propose a new way to study query-key interactions via the Singular Value Decomposition. Many of the modes (i.e. pairs of singular vectors corresponding to the query and the key respectively), are semantic. Two example modes are shown. Top row: ViT layer 8 head 7 mode 2. Bottom row: DINO layer 8 head 9 mode 2. The red channel indicates the projection value of embedding onto the left singular vector which corresponds to the query; the cyan channel indicates the projection value of embedding onto the right singular vector which corresponds to the key.
@@ 1 Introduction
Much like all other deep learning models, though ViTs are successful in many applications, researchers do not have direct access to how information is processed semantically. This issue is particularly important when deploying transformer-based large language models (LLMs) where safety is a priority. As such, there have been studies trying to find feature axes (also known as semantic axes) in the embedding space [16, 4, 5, 12, 31, 17]. A general finding is that embeddings in feedforward layers (i.e. MLP layers) are more semantic than in self-attention layers [16, 17]. It is believed that the embeddings in the self-attention layers have more superposition, whereas embeddings in the feedforward layers have less superposition due to the expansion of dimensionality [5]. Thus, there has been less focus on finding feature axes in the self-attention layers, and there has been little study addressing interactions between feature axes. In this study, while addressing the role of self-attention, we propose that singular vectors of the query-key interaction are pairs of feature directions. Properties of self-attention heads can be elucidated by studying the properties of their singular modes. We show that those singular vector pairs help semantically explain the interaction between tokens in the self-attention layers.
@@
  • We identify a role of self-attention in a variety of ViTs. Specifically, early layers perform more grouping in which tokens attend more to similar tokens; late layers perform more contextualizing in which tokens attend more to dissimilar tokens.
  • We propose a new way to interpret self-attention by analyzing singular modes. Our method goes beyond finding individual feature axes and extends model explainability to the interaction of pairs of feature directions. This approach therefore constitutes progress towards enhancing the explainability of transformer models.
In section 2, we state the motivations of this study and list related work. In section 3, we empirically analyze the preference of self-attention between tokens within and between object categories. In section 4, to study the fundamental properties of the query-key interaction, we propose a Singular Value Decomposition method. In section 5, we show that many of the decomposed singular modes are semantic and can be used to interpret the interaction between tokens. In section 6, we discuss the limitations of this study. In section 7, we discuss the main findings and the significance of this study. In the supplementary, we provide an extensive set of visualization examples of the singular modes.
@@ 2 Related work
The properties of attention maps have been studied since the invention of the ViT. The original ViT paper reported that the model attends to image regions that are semantically meaningful, showing that the $[CLS]$ token attends to objects [13]. Later, a study showed that, in a self-supervised ViT named DINO, the $[CLS]$ attention map has a clearer semantic segmentation property, highlighting the object [6]. Following this idea, studies further showed that the attention map of tokens can highlight parts of an object, and subsequently developed a segmentation algorithm by aggregating attention maps [28, 35]. Other research analyzing the output of self-attention layers indicates that self-attention may perform perceptual grouping of similar visual objects, rather than highlighting a salient singleton object that stands out from other objects in the image [26]. Most of these studies focus on the $[CLS]$ token attention map or on the outputs of attention maps. Our study, in contrast, seeks to interpret the interactions between tokens within the self-attention layers, to gain insights about properties like grouping and contextualization.
@@
Finding feature axes is crucial for understanding and controlling model behavior. Since a study found semanticity in the embeddings of feedforward layers [16], research has primarily focused on finding feature axes in the feedforward layers, and to a lesser extent, in self-attention layers. Bills et al. proposed a gradient-based optimization method to find explainable directions in LLMs [4]. Later, Bricken et al. proposed a simpler method of sparse dictionary learning [5]; though see also [19]. These methods have not been applied to studies of ViTs. However, similar to the findings in LLMs, a ViT study found that feedforward layers have less mixed concepts and can generate interpretable feature visualizations [17].
Some studies focused on finding feature directions in the ViTs’ self-attention layers. In downstream tasks such as semantic segmentation, researchers empirically found that choosing the key embeddings as features leads to the best performance [33, 2, 1]. A study proposed that the singular value decomposition of the weight matrix is a natural way to find feature directions in any neural network [31]. But they only focused on single feature directions (right singular vectors), and did not consider the feature interaction in the context of self-attention. Another study suggested that singular vectors of value weights and feedforward weights can be used as features in LLMs, but they did not analyze the query-key interaction matrix [27].
There has been limited work going beyond single features to studying query-key interactions. A study focusing on LLMs proposed that the corresponding columns of query and key matrices are interpretable as pairs [12]. However, this approach does not find features beyond the standard basis of the query and key embeddings. Here, in contrast to previous works, we utilize the Singular Value Decomposition to study the query-key interactions. We propose that left and right singular vectors of the query-key interaction matrix can be seen as pairs of interacting feature directions, and study their properties in ViTs.
@@ 3 Grouping or contextualizing
Figure 2: Attention preference in the Odd-One-Out (O3) dataset [22]. A. An example from the O3 dataset. Two tokens are chosen to correspond to the target and distractor in the image. Attention maps using two tokens as queries are computed. We examine the overlap between the attention map of the target, and each of the mask labels of the target, distractor, and background masks. Similarly, we examine the overlap between the attention map of the distractor, and each of the mask labels of the distractor, target, and background. B. Ratio of attention on the same objects (target-target and distractor-distractor attention). C. Ratio of attention on the different objects (target-distractor and distractor-target attention). D. Ratio of attention on the background (target-to-background and distractor-background attention)
@@
Firstly, we empirically study whether an image token (i.e. a patch in the image) attends to tokens belonging to the same objects, different objects, or background. We utilized a dataset that has been applied to studying visual salience [22], namely the Odd-One-Out (O3) dataset [26]. This dataset was also used by Mehrami et al [26] in their study but they only focused on the output of the attention layers. However, we use a different experimental design that focuses on the attention maps of image tokens. The dataset consists of 2001 images that have a group of similar objects (distractors) and a distinct singleton object (target) (Fig 2 A). Our goal is to examine if the attention map of a token of one category (target or distractors) covers more of the same category, different category, or background.
@@
We chose to study 12 different ViT models from 4 families: the original ViT [13], DeiT which uses distillation to learn from a teacher model [34], DINO which is trained in a self-supervised way [6], and CLIP which is jointly trained with a text encoder [30].
@@
In this study, the "attention score" is defined as the dot product of every query and key pair, which has the shape of the number of tokens by the number of tokens and is defined per attention head. The "attention map" is the softmax of each query’s attention score reshaped into a 2D image, which is defined per attention head and token. For each image in the dataset, two tokens are chosen to represent the target and distractor. They are at the location of the maximum value of the down-scaled target or distractor mask. Two attention maps are obtained using the two tokens, each is normalized to sum to 1. Inner products are computed between the two attention maps and three masks, which can be interpreted as the ratio of attention of an object (target or distractor) on the same object, different object, or background. We use target-target, target-distractor, target-background, distractor-target, distractor-distractor, and distractor-background attention to denote the 6 inner products. This measure is computed per layer, head, and image. The averaged measure is shown in Fig 2. Target-target and distractor-distractor attention are categorized as "attention on same objects"; target-distractor and distractor-target attention are categorized as "attention on different objects"; target-to-background and distractor-to-background attention are categorized as "attention on background". The attention on the same objects should be dominant if attention is to perform grouping. However, the attention on the same objects is only dominant in early layers. In the deeper layers, there is a trend that attention gradually increases on the contextual features such as the background or different objects. Attention on the backgrounds surpasses that of attention on the same objects in the last few layers.
@@
This result provides new evidence that self-attention considers contextual features as much or more than similar features in deeper layers. The self-attention only prefers the same objects in early layers; in deeper layers, self-attention shifts to prefer contextual information. As far as the authors are aware, this finding has not been reported in previous ViT studies [13, 6, 35, 26].
@@ 4.1 Formulation
We noticed that the attention score solely depends on the combined matrix ${\textbf{W}_{q}}^{\top}\textbf{W}_{k}$ as a whole. To better understand the behavior of this bilinear form, we factor the matrix using the singular value decomposition,
@@ 4.2 Similarity between left and right singular vectors
It is known that embeddings in transformer models are to some extent anisotropic [15, 24, 18], which means the expected value of cosine similarity of two random sampled inputs tends to be positive. We indeed find anisotropy effects in all the models we examined (Supplementary Figure S3). If we treat anisotropy level as a baseline for cosine similarity, the effect shown in Fig 3 still exists but the self-attention is less biased to similar tokens (Supplementary Figure S3).
@@ 5 Semanticity of singular modes
The singular value decomposition of self-attention offers an intuitive way to explain the self-attention layer. A feature represented by a left singular vector attends to the feature represented by the corresponding right singular vector. The feature of a singular vector can be found by finding the image that has the maximum embedding projection on the singular vector. Similarly, the typical interactions of a singular mode can be identified by finding the image that has the maximum product of the projections on a singular vector pair. Previous studies on the explainability of deep learning models only focused on the explainability of single neurons or individual feature axes. The singular value decomposition extends model explainability to the interaction of pairs of "neurons" (i.e. singular vectors).
Some example modes from dino-vitb16 are shown in Fig. 5. For each mode, we choose the top 8 images in the Imagenet (Hugging Face version) [32] validation set that induce the largest attention score. For each image, a query map (red channel in the figure) and a key map (cyan channel in the figure) are obtained by projecting the embedding onto the left and right singular vectors. Each map tells what information the left or right singular vector represents. Jointly, the highlighted regions in the query map attend to the highlighted regions in the key map. In other words, the information in the highlighted regions of the key map flows to the highlighted regions of the query map. More examples are shown for a range of ViT architectures in the Supplementary Figures S4 - S15.
@@
Finally, we study whether tokens prefer to attend to the same object or different objects at the singular mode level. We choose to use a semantic segmentation dataset, namely ADE20K [36]. We first find the top 5 images that induce maximum attention of a singular mode, then find the optimal objects in each image that have the maximum projections on the left and right singular vectors per object area. The probability of the left and right singular vectors having the same optimal object is computed with the weight of singular values, following the same method in the previous experiment. We find that, in early layers, there is a higher probability that the left and right singular vectors attend to the same object; in late layers, the probability is lower, though the variability between models is considerably large. This result further supports that self-attention performs more grouping in early layers; in late layers, tokens attend to different objects which could contextualize the token with background information.

### v3

@@ Abstract
Self-attention in vision transformers is often thought to perform perceptual grouping where tokens attend to other tokens with similar embeddings, which could correspond to semantically similar features of an object. However, attending to dissimilar tokens can be beneficial by providing contextual information. We propose to analyze the query-key interaction by the singular value decomposition of the interaction matrix (i.e. ${\textbf{W}_{q}}^{\top}\textbf{W}_{k}$). We find that in many ViTs, especially those with classification training objectives, early layers attend more to similar tokens, while late layers show increased attention to dissimilar tokens, providing evidence corresponding to perceptual grouping and contextualization, respectively. Many of these interactions between features represented by singular vectors are interpretable and semantic, such as attention between relevant objects, between parts of an object, or between the foreground and background. This offers a novel perspective on interpreting the attention mechanism, which contributes to understanding how transformer models utilize context and salient features when processing images.
Figure 1: We propose a new way to study query-key interactions via the singular value decomposition of the query-key interaction matrix. Many of the modes (i.e. pairs of singular vectors corresponding to the query and the key respectively), are semantically interpretable. Two example modes are shown. Top row: ViT layer 8 head 7 mode 2. Bottom row: DINO layer 8 head 9 mode 2. The red channel indicates the projection value of embedding onto the left singular vector which corresponds to the query; the cyan channel indicates the projection value of embedding onto the right singular vector which corresponds to the key.
@@ 1 Introduction
Much like all other deep learning models, though ViTs are successful in many applications, researchers do not have direct access to how information is processed semantically. This issue is particularly important when deploying transformer-based large language models (LLMs) where safety is a priority. As such, there have been studies trying to find feature axes (also known as semantic axes) in the embedding space [18, 5, 6, 13, 34, 19]. A general finding is that embeddings in feedforward layers (i.e. MLP layers) are more semantically interpretable than in self-attention layers [18, 19]. It is believed that the embeddings in the self-attention layers have more superposition, whereas embeddings in the feedforward layers have less superposition due to the expansion of dimensionality [6]. Thus, there has been less focus on finding feature axes in the self-attention layers, and there has been little study addressing interactions between feature axes. In this study, while addressing the role of self-attention, we propose that singular vectors of the query-key interaction are pairs of feature directions. Properties of self-attention heads can be elucidated by studying the properties of their singular modes. We show that those singular vector pairs help semantically explain the interaction between tokens in the self-attention layers.
@@
  • We identify a role of self-attention in a variety of ViTs. In many ViTs, especially those with classification training objectives, early layers perform more grouping in which tokens attend more to similar tokens; late layers perform more contextualizing in which tokens attend more to dissimilar tokens. However, this observation has some variability among models and may depend on the training objective: notably, some self-supervised ViTs tend to increase attention to dissimilar tokens in the last few layers.
  • We propose a new way to interpret self-attention by analyzing singular modes. Our method goes beyond finding individual feature axes and extends model explainability to the interaction of pairs of feature directions. This approach therefore constitutes enhancing the explainability of transformer models.
In section 2, we state the motivations of this study and list related work. In section 3, we empirically analyze the preference of self-attention between tokens within and between object categories. In section 4, to study the fundamental properties of the query-key interaction, we propose a Singular Value Decomposition method. In section 5, we show that many of the decomposed singular modes are semantic and can be used to interpret the interaction between tokens. In section 6, we discuss the limitations of this study. In section 7, we discuss the main findings and the significance of this study. In the supplementary, we provide an extensive set of visualization examples of the singular modes. Code for this work is available at: https://github.com/schwartz-cnl/DissectingViT.
@@ 2 Related work
The properties of attention maps have been studied since the invention of the ViT. The original ViT paper reported that the model attends to image regions that are semantically meaningful, showing that the $[CLS]$ token (i.e. a special token originally designed as the final hidden vector) attends to objects [14]. Later, a study showed that, in a self-supervised ViT named DINO, the $[CLS]$ attention map has a clearer semantic segmentation property, highlighting the object [7]. Following this idea, studies further showed that the attention map of tokens can highlight parts of an object, and subsequently developed a segmentation algorithm by aggregating attention maps [31, 39]. Another study on the output of self-attention layers indicates that self-attention may perform perceptual grouping of similar visual objects, rather than highlighting a salient singleton object that stands out from other objects in the image [29]. Most of these studies focus on the $[CLS]$ token attention map or on the outputs of attention maps. Our study, in contrast, seeks to interpret the interactions between tokens within the self-attention layers, to gain insights about properties such as grouping and contextualization.
@@
Finding feature axes is crucial for understanding and controlling model behavior. Since a study found semanticity in the embeddings of feedforward layers in LLMs [18], studies have primarily focused on identifying feature axes in the feedforward layers, and to a lesser extent, in self-attention layers. Similar to the findings in LLMs, a ViT study found that feedforward layers have less mixed concepts and can generate interpretable feature visualizations [19]. Bills et al. proposed a gradient-based optimization method to find explainable directions in LLMs [5]. Later, Bricken et al. proposed a simpler method of sparse autoencoder [6]; though see [21]. These methods have not been extensively applied to ViT studies.
Some studies focused on finding feature directions in the ViTs’ self-attention layers. In downstream tasks such as semantic segmentation, researchers empirically found that choosing the key embeddings as features leads to the best performance [37, 2, 1]. A study proposed that the singular value decomposition of the weight matrix is a natural way to find feature directions in any neural network [34]. But they only focused on single feature directions (right singular vectors), and did not consider the feature interaction in the context of self-attention. Another study suggested that singular vectors of value weights and feedforward weights can be used as features in LLMs, but they did not analyze the query-key interaction matrix [30]. Another study in the language domain proposed a singular vector decomposition on the union of the query and key embeddings, but not on the query and key weights [27].
There has been limited work going beyond single features to studying query-key interactions. A study focusing on LLMs proposed that the corresponding columns of query and key matrices are interpretable as pairs [13]. However, this approach does not find features beyond the canonical basis of the query and key embeddings. Another study inferred query-key interactions by jointly visualizing them in a low dimensional space, but their method does not find interacting feature axes [42]. Here, in contrast to previous works, we utilize the singular value decomposition to study the query-key interactions. We propose that left and right singular vectors of the query-key interaction matrix can be seen as pairs of interacting feature directions, and study their properties in ViTs.
@@ 3 Grouping or contextualizing
Figure 2: Attention preference in the Odd-One-Out (O3) dataset [24]. A. An example from the O3 dataset. Two tokens are chosen to correspond to the target and distractor in the image. Attention maps using two tokens as queries are computed. We examine the overlap between the attention map of the target, and each of the mask labels of the target, distractor, and background masks. Similarly, we examine the overlap between the attention map of the distractor, and each of the mask labels of the distractor, target, and background. B. Ratio of attention on the same objects (target-target and distractor-distractor attention). The x-axis is normalized layer numbers, from early layers (left) to late layers (right). C. Ratio of attention on the different objects (target-distractor and distractor-target attention). D. Ratio of attention on the background (target-to-background and distractor-background attention)
@@
Firstly, we empirically study whether an image token (i.e. a patch in the image) attends to tokens belonging to the same objects, different objects, or background. We utilized a dataset that has been applied to studying visual salience [24], namely the Odd-One-Out (O3) dataset [29]. This dataset was also used by Mehrami et al. [29] in their study but they only focused on the output of the attention layers. However, we use a different experimental design that focuses on the attention maps of image tokens. The dataset consists of 2001 images that have a group of similar objects (distractors) and a distinct singleton object (target) (Fig 2 A). Our goal is to examine if the attention map of a token of one category (target or distractors) covers more of the same category, different category, or background.
@@
We chose to study 16 different ViT models from 6 families: the original ViT [14], DeiT which uses distillation to learn from a teacher model [38], CLIP which is jointly trained with a text encoder [33], and DINO [7], SimMIM [40], I-JEPA [4] which are self-supervised models with either contrastive or mask prediction loss.
@@
In this study, the "attention score" is defined as the dot product of every query and key pair, which has the shape of the number of tokens by the number of tokens and is defined per attention head. The "attention map" is the softmax of each query’s attention score reshaped into a 2D image, which is defined per attention head and token. For each image in the dataset, two tokens are chosen to represent the target and distractor. They are at the location of the maximum value of the down-scaled target or distractor mask. Two attention maps are obtained using the two tokens, each is normalized to sum to 1. Inner products are computed between the two attention maps and three masks, which can be interpreted as the ratio of attention of an object (target or distractor) on the same object, different object, or background. We use target-target, target-distractor, target-background, distractor-target, distractor-distractor, and distractor-background attention to denote the 6 inner products. This measure is computed per layer, head, and image. The averaged measure is shown in Fig 2. Target-target and distractor-distractor attention are categorized as "attention on same objects"; target-distractor and distractor-target attention are categorized as "attention on different objects"; target-to-background and distractor-to-background attention are categorized as "attention on background". The attention on the same objects should be dominant if attention is to perform grouping. We find a trend that in most ViTs the attention on the same objects is dominant in early layers; while there is a trend that in the deeper layers attention gradually increases on the contextual features such as the background or different objects. However, this observation has some variability among models and may depend on the training objective. For example, the self-supervised SimMIM pre-trained on pixel-level mask prediction shows increased attention on the same objects in later layers. Interestingly, this trend disappears after fine-tuning on a classification task.
@@
This result provides new evidence that self-attention considers contextual features as much or more than similar features in deeper layers. In most ViTs, especially with those with classification training objectives, self-attention prefers the same objects in early layers; in deeper layers, self-attention shifts to contextual information. As far as the authors are aware, this finding has not been reported in previous ViT studies [14, 7, 39, 29].
@@ 4.1 Formulation
The attention score solely depends on the combined matrix ${\textbf{W}_{q}}^{\top}\textbf{W}_{k}$ as a whole [15], which represents the query-key interaction. To better understand the behavior of this bilinear form, we factor the matrix using the singular value decomposition,
@@ 4.2 Similarity between left and right singular vectors
Though we find a general trend that attention changes from attending more to the similar tokens to dissimilar tokens from early layers to late layers, some ViTs have a more complex trend that increases attention to similar tokens in the last few layers (Fig 3). Models that have this “concave” trend are SimMIM-vit-b16-pretrain, Dino models, Deit models, and huge ViT models. Most of them either have self-supervised objectives or distillation regularizations. We hypothesize that the last layers may behave differently because they are closer to the training target, and so the training objective may have more influence. We think that self-supervised objectives, such as reconstructing masked patches, require stronger consistency between tokens, and thus more attention is allocated to similar tokens in the higher layers; while the classification objective requires gathering information from different aspects of a scene, and thus more attention is allocated to dissimilar tokens. This hypothesis is supported by the cosine similarity plot (Fig 3) of the SimMIM models, which shows in the last few layers of the pre-trained model increased attention to similar features. This matches the observation in the literature, that the SimMIM model has more local attention [41]. However, we find that the SimMIM model fine-tuned on ImageNet classification has a trend of decreased attention to similar features, similar to most of the classification models. Although I-JEPA is trained with a self-supervised objective predicting latent representations, the cosine similarity for the I-JEPA encoder does not show increased attention to similar tokens in the last few layers. The I-JEPA model is known to have excellent linear-probing performance, and thus we think it may behave more similarly to a classification model. The self-supervised objective of I-JEPA may be more apparent in the I-JEPA predictor (also a transformer). When we run the cosine similarity analysis on the predictor module instead of the encoder, we find that the cosine similarity is overall high (Supplementary FigureS3). The role of the training objective on internal model behavior is an interesting topic for future research.
It is known that embeddings in transformer models are to some extent anisotropic [17, 26, 20], which means the expected value of cosine similarity of two random sampled inputs tends to be positive. We indeed find anisotropy effects in all the models we examined using cosine similarity (Supplementary Figure S4) (though see other metrics [35]). If we treat anisotropy level as a baseline for cosine similarity, the effect shown in Fig 3 still exists but the self-attention is less biased to similar tokens (Supplementary Figure S4).
@@ 5 Semanticity of singular modes
Figure 5: Visualization of a single image with multiple modes. We pick an example dog image from the ImageNet dataset and use the dino-vitb16 model. Top 6 modes (ordered by the contribution to the attention score) for example layers and heads are shown. See Supplementary Figure S17 for extended mode visualizations of this image.
The singular value decomposition of self-attention offers an intuitive way to explain the self-attention layer. A feature represented by a left singular vector attends to the feature represented by the corresponding right singular vector. The feature of a singular vector can be found by finding the image that has the maximum embedding projection on the singular vector. Similarly, the typical interactions of a singular mode can be identified by finding the image that has the maximum product of the projections on a singular vector pair. Previous studies on the explainability of deep learning models only focused on the explainability of single neurons or individual feature axes. The singular value decomposition extends model explainability to the interaction of pairs of "neurons" (i.e. singular vectors). Note that this is very different from the standard approach of visualizing the attention map of the $[CLS]$ token without addressing interactions between tokens [14, 7, 31].
Some example modes from dino-vitb16 are shown in Fig. 4. For each mode, we show the top 8 images in the Imagenet (Hugging Face version) [36] validation set that induce the largest attention score. For each image, a query map (red channel in the figure) and a key map (cyan channel in the figure) are obtained by projecting the embedding onto the left and right singular vectors. Each map tells what information the left or right singular vector represents. Jointly, the highlighted regions in the query map attend to the highlighted regions in the key map. In other words, the information in the highlighted regions of the key map flows to the highlighted regions of the query map. More examples are shown for a range of ViT architectures in the Supplementary Figures S5 - S16.
@@
To show the hierarchical information process across layers, we show an example dog image and example attention heads along with optimal images for top modes in Fig. 5. The late layers capture more semantic information such as the parts of a dog or animal, and a hand with a dog. The early layers capture low-level properties like color. We show more examples in Supplementary Figure S17-S19.
@@
Finally, we study whether tokens prefer to attend to the same object or different objects at the singular mode level. We choose to use a semantic segmentation dataset, namely ADE20K [43]. We first find the top 5 images that induce maximum attention of a singular mode, then find the optimal objects in each image that have the maximum projections on the left and right singular vectors per object area. The probability of the left and right singular vectors having the same optimal object is computed with the weight of singular values, following the same method in the previous experiment. We find that, in early layers, there is a higher probability that the left and right singular vectors attend to the same object; in late layers, the probability is lower, though the variability between models is considerably large (Fig. 6). This result further supports that self-attention performs more grouping in early layers; in late layers, tokens attend to different objects which could contextualize the token with background information.